FEL Upgrade Project Weekly Brief – September 16-20, 2002

We've now established the required field uniformity of +/- 2 parts in 10000 in the GV magnets (used in the injector chicane) by the addition of small shunts.  The establishment of this approach and the hysteresis protocol sets the stage for final adjustment and measurement of these magnets prior to release to installation.

We measured the HOM damping on a 7-cell cavity with the HOM couplers and the high power couplers attached.  On the basis of this measurement we have given the go-ahead to proceed to fabrication of the FEL cavities for the new style cryomodule to be installed next summer.

We powered up one klystron and put rf in the (shorted) injector waveguide to the tunnel for the first time.

We are assembling the cost data and earned value for August so we can release the monthly report at the beginning of next week.

H. F. Dylla and R. Rimmer from the SRF group participated in a review of the Duke FEL program this week.
M. Shinn attended the Boulder Damage Conference and presented a paper on our FEL activities.

We reviewed the optical system construction progress against the master schedule and appear to be on track for producing all systems required for initiation of lasing by the time the electron beam will be commissioned and ready.

As major hardware is getting installed and turning things on gets close, coordination efforts are becoming more intense but so far no major roadblocks have occurred.  We have approved overtime work in a few of key areas.  We are still on track for start of commissioning at the end of the month as originally planned (not to say there isn't a lot yet to do....).

WBS 4 (Injector):
The rear bellows/drive unit and rear corona shield were submitted to the shop for fabrication. The gun stand was grouted in place and set up/cleaning of the of the clean enclosure around the gun chamber was done.  Assembly of the electrode structure into the ceramic stack was started in the off line clean room of the test lab.  Assembly of the light box revealed some off sets which are being corrected by re-machining of parts.  Field measurements of the guns reworked solenoid were started.

Gun HVPS - The tank flanges are being wired and will be installed next week.  We will be able to start testing the HVPS late next week.  The Interlocks have been determined and are being wired.  We still need a larger or an additional gas bag for the SF6 when the tank is pumped out and opened.

WBS 5 (SRF):
We summarized the measurements of EP3 in the vertical test and of the copper model with additional elliptical HOM probe on the FPC end.  The FPC termination seems to produce a useful improvement in one orientation of the TE111#5 mode at 1871 MHz but otherwise no major differences with previous HOM-only data.  This should at least raise the threshold of one mode in the FEL.  The elliptical HOM probe added to the FPC end of the copper model also produced a useful improvement in one bad mode, #11 in Kevin's table, raising the threshold by almost a factor of 10, and improved mode #13 by about a factor of two.  Unfortunately it had no real effect on mode #5, at least in these measurements. In some cases the Q's are higher than EP3 despite the added HOM antenna, which may be due to the lack of field flatness of the copper model.  Inspection of the SL21 string shows how little room there is to make any changes to the end groups.

Given the results to date and the technical difficulty of moving the couplers closer and the potential schedule impact of delays trying to improvedamping further we have decided to proceed with construction of the cavity as presently designed.  We will continue to study the damping for possible improvements in future modules.

In fabrication, only two FEL cavities have been welded, which is the same as last week.  The SL21 string continues to gain components. Tuners have been fitted up and removed again for modifications. Magnetic shields are in place around the He vessels

WBS 6 (RF):
Quarter HVPS - RF testing started this afternoon on Unit 3 under Local Control.  The LCW was connected the Circulator, RF Load, and HOM Filter this week.  The software is making good progress and is able to communicate with the hardware.  Additional software testing will be done on Unit 3 next week while hardware installation and checkout continues on Unit 4.

Quarter Klystrons - Work continues on klystron carts 2 & 3.

WBS 8 (Instrumentation):
The vacuum rack is nearing completion, currently the ion pump power supplies are downstairs local to the pumps.  Next week the 3 racks will be recommissioned as well as the new software and all the controls and readout will be available in EPICS.

The new MPS boards have gone out for fabrication.  These will eliminate the need for any masking of hardware devices in the upgrade.  We began the MPS rack rework, this eliminates the patch cables, PLCs, and terminal blocks.  The design was completed for the control room interface box.

The dump water skid was refilled and recommissioned this week.  One of the conductivity monitors had failed and is being replaced, both of the sensors were replace as preventative maintenance.  Railings and crossovers were installed in the 3F region to prevent any trip accidents for people trying to get over the beamline.

Work is progressing on the HV interlock chassis with completion and installation to follow next week.  The Acromag fanout module has been finished and sent out to a board house with a turnaround date of 5 days.  A meeting was held on Tuesday with the respective folks from the Software, Operations and FEL groups to format the new trim magnet user screens for the 0F region.  The Trim Magnet software is well in progress. 40 Beamviewer Interface boxes have been recieved from the machine shop and parts for the boards have been placed on order.

BLMs - The F0150 rev A (rear cards) are back from the board house and look great. We immediately assembled one to serve for the first 12 BLM channels to come online by Oct. 1st. The remaining boards will be implemented later in the commissioning process as priority demands.  The F0151 rev - prototype is fully populated and mostly programmed.  The card set is currently installed into and communicating with iocfelt2 in Lab 5 where it will finish out its testing next week and then be installed into the FEL proper. DLPC - The ConOptics carrier board is out for production. In the mean-time, work has begun to implement the automatic RF switch for the internal vs. external RF for the drive laser. MISC - The photocathode scanner chassis has been recommissioned in Lab 5 and was then given to the Optics group for in-system testing.  After a successful functionality test, the chassis is back with us to be further documented.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
•  MEI is checked all the core gaps of the remaining six cores and found that the upper and lower
    core pieces were flat but that the return legs were not.  They are having the return legs kissed at
    the grinders to bring them into tolerance.
•  Magnet Test will activate the field integral match stand on two GW magnets on Monday to start
    their formal testing
Injector Dipoles (DU/DV)
•  By using various shims on the backside of the field clamp edges on either side of the center, we
    were able to bring the integral linearity (It had sextupole) of the GV to within 2 parts in 10,000.
    (Observed with a course scan.)  When we used the fine scan, the integral linearity is 4 parts in
    10,000 and we can see that we have a very powerful tool to adjust the field in to specification
    with the shims.  At week’s end, we had modified field clamps to use bolted support plates under
    the shims (rather than double sided tape that allowed the shims to peal off) and we are homing in
    on the shim stack thickness and position that will bring the integral linearity well within specification
    for all magnets of the series.
Arc 180 Degree Dipoles (GY)
•  GY Coil Potting fixture continues its preparation for potting.
•  The extra copper for the third and forth layers of the 5th coil was received by Wang NMR.  They
    had completed winding the first two layers earlier in the week.
•  Wang started preliminary work on gluing the shims down on the GY cores.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
•  At Wang NMR, the second GQ coil was vacuum impregnated with epoxy and first coil passed
    hi-pot test.  They believe they can ship the first magnet next Friday.
Trim Quad (QT)
•  The second trim quad was fully measured.
•  They found that the field from the trim quadrupole test affected the field integral testing on the
    stepper stand such that the tests had to be done only when the stepper stand was not working.
Sextupole (SF)
•  DULY Research passed all the magnet drawings to us.
•  A preliminary layout indicates that the quick fix sextupoles can be made from the IR Demo
    sextupoles and the simple addition of some shim pieces, some altered coil supports and some
    redrilling of the original core pieces.
Octupole (OT)
•  The Octupole design continued at DULY Research where they generated their first preliminary
    core and pole tip drawings.
Beam Line and Vacuum
•  X-Chamber - Welding of transitions to chambers continues to be a challenge.  We needed to
    remove material in order to obtain a full penetration joint.  The joints are being re-welded.  The
    experience learned from this effort will be used to modify the weld prep on the other chamber
    transition interfaces.
•  Chicane Chambers - Welding on the first chicane chamber continued while X-chamber was on the
    mill.  Wire EDM work on the chicane transition is in process.  Machining of other chamber plates,
    tubes, and CFF's are on hold to allow higher priority parts to be fabricated.
•  Y Chamber - Plate machining is in progress and is almost complete.
•  ARC Chambers - Master Machine continues work on the first ARC chamber.  Full penetration
    fusion weld is progressing (as inspected by Bore scope.  Some difficulty was experienced at the
    transition to chamber interface (similar to the problems experience at JLab shop).  A similar
    rework effort is producing acceptable results.
The machining of the bevels is still delayed due to a controller problem on the CNC machine.  The machine has been repaired and is being run through some initial testing.  Return to service is expected soon.  A subcontractor has been utilized to keep machining efforts moving.
•  We signed off the location drawing for the trim quad and sextupole stands.
•  Girder and vacuum assembly continued
•  Low Conductivity Water piping installation continued in the FEL enclosure, it is now more than
    50% complete.
•  A number of stands are ready for a grouting party.

WBS 11 (Optics):
The leadscrews for the optical cavities have been sent to Thin Films, Inc, to be sputtered with tungsten disulfide.  We received all the LVDTS, prepared for high vacuum use, required for the optical cavity and optical transport to the OCR.  Outgassing checks are ongoing at the time of this report.  The OCMMS feedback control is progressing, with dampening by about a factor of 2 in the 1-15 Hz region, in a very noisy environment.  Parts of the optical transport hardware are being procured.
We installed a HeNe in the drive laser clean room in order to align its optical transport.  The transport is aligned to the last turning mirror over the light box.  Some work remains to insure everything is firmly attached before the relay telescope is installed.  The optical components for the photocathode scanner were modified for the new light box and are ready for installation.

Other activities:
Michelle Shinn gave an invited talk at the 34th Annual Symposium on Optical Materials for High Power Lasers.  While there, she learned the results of the latest research on the optical properties of calcium fluoride (useful for the UV FEL), and on laser beam characterization.  She also met with several optics vendors.